If the length of a rectangular parking lot is 10 tard less than twice its width. if the perimeter of the lot is 400 yrads, what are its dimention?
Answer by xcentaur(357)

let the length be x and width be y
then
x=2y-10.......[1]
2(x+y)=400....[2]


substituting [1] in [2] we get
2(2y-10+y)=400
(2y-10+y)=200
3y-10=200
3y=210
y=70


x=2y-10=140-10=130


Dimensions are: 70 yards,130 yards


Hope this helps,
Prabhat